# Towards the New XAI: A Hypothesis-Driven Approach to Decision Support Using Evidence

### v3

@@ 3 Evidence-Informed Hypothesis-Driven Decision-Making Model
In a classification problem, a hypothesis $h\in Y$, where $Y=\{h,h_{1},h_{2},...,h_{n}\}$ includes all possible hypotheses, as an output class. Then, $\bar{h}=Y\setminus{\{h\}}$ refers to all hypotheses other than $h$. For example, if a doctor asserts a set of hypotheses $Y=\{h_{1},h_{2},h_{3}\}$ where $h_{1}$ = the patient has Covid, $h_{2}$ = the patient has Influenza and $h_{3}$ = the patient has pneumonia, then $\bar{h_{1}}$ = the patient does not have Covid which includes all possible hypotheses except having Covid, that is $\bar{h_{1}}=\{h_{2},h_{3}\}$.
@@
We generate the weight of evidence for possible hypotheses using Weight of evidence (WoE), which is a probabilistic approach for analysing variable importance, introduced in the context of explainability by Alvarez Melis et al. [2] building on the approach of Good [17]. It provides a quantitative response to the question of why a model predicted output $h$ for a particular input $x$ in terms of how much each input feature $x_{i}$ provides in favour of, or against, $h$, relative to alternatives. Through Bayes rule, WoE can be understood as an adjustment to the prior log odds caused by observing the evidence.
@@
$\operatorname*{woe}(h\mid x_{i})=\log\frac{P(x_{i}\mid h)}{P(x_{i}\mid\bar{h})}=\log\frac{P(h\mid x_{i})}{P(\bar{h}\mid x_{i})}-\log\frac{P(h)}{P(\bar{h})}$ (1)
@@
How decision-aid models can use WoE to make a decision
@@
How WoE can incorporate a human approach to making a decision
To assist users with interpretability, Alvarez Melis et al. [2] complement the display of the magnitude of the weight of the evidence with a notion of significance level of the evidence, using a scale of seven categories: decisive-against (), strong-against (), substantial-against (-), not-significant (N), substantial-in-favor (+), strong-in-favor (++), decisive-in-favor ($+$). The details can be found in the rule-of-thumb guidelines here [3].
In addition to the weight of evidence of a feature, we suggest it is useful to distinguish the importance of a feature – with importance being domain specific and determined by the domain expert using the model. Specifically, if a feature has significant weight of evidence according to the WoE model, but that feature is not seen as important by the human decision maker, then it is reasonable to anticipate the impact of that evidence on the decision would be reduced by the decision maker. For example, if a clinician looks at a dermatoscopic image and also is aware of some irrelevant but high weight of evidence feature such as dense hair, they should ignore that evidence in making a prediction.
@@
where $\gamma_{i}$ is a parameter of feature $x_{i}$ that adjusts the weight of evidence based on importance, i.e., $\gamma_{i}>\gamma_{j}$ represents that feature $x_{i}$ is more important than feature $x_{j}$.
@@
4 Experiment Design
In this section, we describe the task implemented in the human behaviour experiment and the experiment design. In selecting a decision-making task, we identified requirements similar to those used in other studies of how explanations can assist human decision makers interacting with AI decision support, e.g. Vasconcelos et al. [37]: the task should not be too easy for humans to complete without a decision aid, but also, as we were using lay subjects from Prolific for this particular study, the task cannot require specialist knowledge.
We chose a version of the housing price prediction task studied previously in an XAI context [1, 10]. In this task, participants are provided with information about house features and with other information which varies by experimental condition, and are asked to choose whether the given house would have a sale price of low, medium or high. As noted by others [10], real estate valuation is a domain where ML models have been developed to help people make better decisions, predicting house prices is a task that lay people may need to do in real life, so it is not unrealistic to expect they have sufficient day-to-day knowledge to make predictions and decide whether or not to rely on an AI model.
@@ 4 Experiment Design
Experimenting with this task, we compared the hypothesis-driven approach with two state-of-the-art decision-making approaches using quantitative measures for efficiency, performance and reliance and a qualitative analysis of information use. In the terminology of a recent review of XAI evaluation [22], the first two points of comparison are a form of evaluation with respect to the decision task, and the latter two focus on users’ perception and use of the AI system itself.
@@ 4.1 Dataset and Model Implementation
We selected six features for the human experiment in the house-price decision-making task by applying Gradient Boosting Classification model over the data. Considering domain specific decision-making about house prices, we propose there to be three important features (quality of construction, house age and location) and three unimportant features (fireplaces, kitchen quality and central air conditioning). We divided the dataset into 80% for the training set and 20% for the test set. Following Alvarez Melis et al. [2], we use a Gaussian Naïve Bayes (GNB) classifier to obtain $P(x_{i}\mid h)$. This assumes that features are independent, but the model and implementation work for any probabilistic classifier. We chose this model because it is a simple discriminative classifier that aligns with previous work on evidence-based explanations [29, 21].
@@ 4.2 Experimental Conditions
All participants [^1] We received ethics approval from our institution before conducting the human experiment. were given the six house feature values plus other information, which varied by condition as set out below. Participants then chose whether the given house would have a price of low, medium or high. Using a between-subject design, participants were randomly assigned to one of three conditions:
@@ 4.3 Research Questions and Hypotheses
  • RQ4: (Information use) How do people make decisions differently in recommendation-driven, AI-explanation-only and hypothesis-driven paradigm?
@@
  • H1a/b: (C3) Hypothesis-driven paradigm will cost less time to finish the task than (C1) Recommendation-driven and (C2) AI-explanation-only.
@@
  • H2a/b: (C3) Hypothesis-driven paradigm will help participants make better decisions than (C1) Recommendation-driven and (C2) AI-explanation-only.
@@
For RQ4, we looked into the text written by participants when they explained why they selected an option after each question to know how they used the provided information in each decision-making paradigm to make their decisions. Therefore, we can identify the limitations of each paradigm and the generated evidence that lead the participants to make a wrong decision.
@@ 4.4 Measures
    where: $C_{p,i}$ is the likelihood level of participant $p$ in question $i$, ranging from 0 to 1; $A_{p,i}$ is the answer score of participant $p$ in question $i$, either 0 (wrong answer) or 1 (right answer); N is the number of questions for each participant. The best Brier score (i.e. equal to 0) for an individual task is when a participant answers the task correctly and gives it a 100% likelihood (or alternatively, a wrong answer but with 0% likelihoood). Therefore, a participant has better task performance when they have a lower Brier score. The Brier score measures decision accuracy, but awards a higher score for a correct answer when a participant is confident, and lower penalty for an incorrect answer when not confident. This mitigates problems where participants guess answers (i.e. have low confidence in their answers).

### v4

@@ 3.1 Weight of Evidence
In a classification problem, a hypothesis $h\in Y$, where $Y$ is a set of possible hypotheses. Then, $Y_{-h}=Y\setminus{\{h\}}$ refers to all hypotheses other than $h$. For example, if a doctor asserts a set of hypotheses $Y=\{h_{1},h_{2},h_{3}\}$ where $h_{1}$ = the patient has Covid, $h_{2}$ = the patient has Influenza and $h_{3}$ = the patient has pneumonia, then $Y_{-h_{1}}$ = the patient does not have Covid which includes all possible hypotheses except having Covid, that is $Y_{-h_{1}}=\{h_{2},h_{3}\}$.
@@
We generate the weight of evidence for possible hypotheses by applying the Weight of Evidence (WoE) framework, which is a probabilistic approach for analysing variable importance, introduced in the context of explainability by Melis et al. [59] building on the approach of Good [26]. It provides a quantitative response to the question of why a model predicted output $h$ for a particular input $X$ in terms of how much each input feature $x_{i}$ provides in favour of, or against, $h$, relative to alternatives. Through Bayes rule, WoE can be understood as an adjustment to the prior log odds caused by observing the evidence.
@@
$\operatorname*{woe}(h\mid x_{i})=\log\frac{P(x_{i}\mid h)}{P(x_{i}\mid Y_{-h})}=\log P(x_{i}\mid h)-\log P(x_{i}\mid Y_{-h})$ (1)
where $\log P(x_{i}\mid h)$ is the Gaussian log density for hypothesis $h$.
@@
Considering the vector input $X=[x_{1},x_{2},\dots,x_{n}]$, the features can be independent or dependent on each other. We can apply Gaussian log density for both independent and dependent variables, depending on how we handle the covariance matrix $\Sigma$. The covariance matrix measures the relationship between two variables, indicating the direction (positive or negative) of how much each pair of features changes together and the strength of their relationship. For independent variables, the covariance matrix is diagonal, while for dependent variables, the covariance matrix is full.
Independent Variables
For independent variables, the Gaussian log density is based on a univariate Gaussian distribution, which can be simplified as follows:
$P(x_{i}|h)$ $=\mathcal{N}(x_{i};\mu_{i|h},\sigma_{i|h}^{2})=\frac{1}{\sqrt{2\pi\sigma_{i|h}^{2}}}\exp\left(-\frac{(x_{i}-\mu_{i|h})^{2}}{2\sigma_{i|h}^{2}}\right)$
$\log P(x_{i}|h)$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{2}\log(\sigma_{i|h}^{2})-\frac{(x_{i}-\mu_{i|h})^{2}}{2\sigma_{i|h}^{2}}$ (2)
where $\mu_{i,h}$ is the mean of $x_{i}$ for hypothesis $h$, and $\sigma_{i,h}^{2}$ is the variance of $x_{i}$ for hypothesis $h$.
Dependent Variables
For dependent variables, the Gaussian log density is based on a multivariate Gaussian distribution. In this case, we compute the conditional distribution of $x_{i}$ given the remaining features $X_{-i}$ and the hypothesis $h$. We denote i-i as the set of features excluding the feature $i$. We calculate conditional mean, conditional variance and Gaussian log density as follows.
$\mu_{i|-i,h}$ $=\mu_{i|h}+\Sigma_{i,-i|h}\Sigma_{-i,-i|h}^{-1}(X_{-i}-\mu_{-i|h})$
$\sigma_{i|-i,h}^{2}$ $=\Sigma_{i,i|h}-\Sigma_{i,-i|h}\Sigma_{-i,-i|h}^{-1}\Sigma_{-i,i|h}$
$P(x_{i}|X_{-i},h)$ $=\mathcal{N}(x_{i};\mu_{i|-i,h},\sigma_{i|-i,h}^{2})=\frac{1}{\sqrt{2\pi\sigma_{i|-i,h}^{2}}}\exp\left(-\frac{(x_{i}-\mu_{i|-i,h})^{2}}{2\sigma_{i|-i,h}^{2}}\right)$
$\log P(x_{i}|X_{-i},h)$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{2}\log(\sigma_{i|-i,h}^{2})-\frac{(x_{i}-\mu_{i|-i,h})^{2}}{2\sigma_{i|-i,h}^{2}}$ (3)
Subset of features
When we consider a subset of features $S$ for hypothesis $h$, we can calculate the conditional distribution of $X_{S}$ given the remaining features $X_{-S}$ and the hypothesis $h$. We denote S-S as the set of features excluding $S$. We calculate the conditional mean, conditional variance matrix and Gaussian log density as follows.
$\mu_{S|-S,h}$ $=\mu_{S|h}+\Sigma_{S,-S|h}\Sigma_{-S,-S|h}^{-1}(X_{-S|h}-\mu_{-S|h})$
$\Sigma_{S|-S,h}$ $=\Sigma_{S,S|h}-\Sigma_{S,-S|h}\Sigma_{-S,-S|h}^{-1}\Sigma_{-S,S|h}$
$\log P(x_{S}|x_{-S},y)$ $=-\frac{|S|}{2}\log(2\pi)-\frac{1}{2}\log|\Sigma_{S|x_{-S},y}|-\frac{1}{2}(x_{S}-\mu_{S|x_{-S},y})^{\top}\Sigma_{S|x_{-S},y}^{-1}(x_{S}-\mu_{S|x_{-S},y})$ (4)
Mixture model
To compute $P(x_{i}|Y_{-h})$, we sum over all possible hypotheses $k\in Y_{-h}$:
$P(x_{i}|Y_{-h})=\sum_{k\in Y\setminus{\{h\}}}P(x_{i}|k)P(k)$ (5)
3.1.1 How decision-aid models can use WoE to make a decision
@@
3.1.2 How WoE can incorporate a human approach to making a decision
To assist users with interpretability, Melis et al. [59] complement the display of the magnitude of the weight of the evidence with a notion of the significance level of the evidence, using a scale of seven categories: decisive-against (), strong-against (), substantial-against (-), not-significant (N), substantial-in-favour (+), strong-in-favour (++), decisive-in-favour (+++). The details can be found in the rule-of-thumb guidelines here [2].
In addition to the weight of evidence of a feature, we suggest it is useful to distinguish the importance of a feature – with importance being domain-specific and determined by the domain expert using the model. Specifically, if a feature has a significant weight of evidence according to the WoE model, but that feature is not seen as important by the human decision-maker, then it is reasonable to anticipate the impact of that evidence on the decision would be reduced by the decision maker. For example, if a clinician looks at a dermatoscopic image and also is aware of some irrelevant but high weight of evidence features such as dense hair, they should ignore that evidence in making a prediction.
@@ 3.1.2 How WoE can incorporate a human approach to making a decision
where $\gamma_{i}$ is a parameter of feature $x_{i}$ that adjusts the weight of evidence based on importance, i.e., $\gamma_{i}>\gamma_{j}$ represents that feature $x_{i}$ is more important than feature $x_{j}$. $\gamma$ can be considered as prior belief of the human decision-maker.
@@
3.2 Visual Evaluative AI
The Weight of Evidence (WoE) framework in Section 3.1 can be applied to tabular datasets. However, extracting features from images is more complicated than that as it requires sophisticated techniques like convolutional neural networks. Therefore, we extend the WoE by applying concept-based explanations [88, 89] to extract human-understandable concepts from images and put these concepts into the WoE model. Additionally, despite existing methods like Grad-CAM [73] can provide contrastive explanations of various output classes, pixel-level explanations are less effective as they only specify which area of the image is important, but not what is important about that area. [35].
In this section, we introduce our evidence generation model by combining a concept-based explanation model (e.g., Invertible Concept-based Explanation (ICE) [89], Post-hoc Concept Bottleneck Model (PCBM) [88]) and the Weight of Evidence (WoE) model [59]. In our experiments, we use Invertible Concept-based Explanation (ICE) [89] as an example of unsupervised concept learning, and Post-hoc Concept Bottleneck Model (PCBM) [88] as an example of supervised concept learning. Combining them together, we propose two models to generate the evidence-based explanations called ICE+WoE and PCBM+WoE.
3.2.1 Concept-based Explanations
We divide concept-based explanations into two categories: (1) supervised learning concepts (concepts are labelled on each image in the training dataset) and (2) unsupervised learning concepts (not having concept labels in the training dataset). Supervised concept learning requires labelled concepts in the training set, or the concepts can be transferred using another labelled dataset [88]. Unsupervised learning concept methods do not require the concepts to be labelled during the training process. This method is helpful when labelling concepts can be laborious, require expertise, or are not always available. Moreover, unsupervised learning can give users more agency as they can find a new concept that has not been labelled but is still used by a machine learning model. In our application, the evidence will be referred to a concept (or feature) found in the image. Each concept will have a positive/negative quantitative value that shows how much it contributes to the given hypothesis.
The main difference between ICE+WoE and PCBM+WoE is that the concepts generated by ICE+WoE do not have labels returned by the model. In other words, ICE+WoE identifies only important concepts for the classifier but does not assign a name for them. The unlabelled concepts can then be assigned labels by a domain expert. On the other hand, PCBM+WoE provides a concept name for each concept, which is learned from the concept bank. It is important to note that the labelled concepts are not always reliable and still require validation by a domain expert. Furthermore, the number of concepts is fixed based on the concept bank in PCBM+WoE, while the number of concepts can be chosen by the user in ICE+WoE.
Figure 1: Unsupervised Concept Learning Model (ICE and ICE+WoE)
Figure 2: Supervised Concept Learning Model (PCBM and PCBM+WoE)
3.2.2 ICE+WoE and PCBM+WoE
We will now explain the implementation of the combined models of concept-based explanations and the Weight of Evidence (WoE) model, depending on the concept learning method (unsupervised or supervised). Figure 1 shows an overview of the unsupervised concept learning model, and how it can be combined with Weight of Evidence (WoE) [59]. Figure 2 shows an overview of the supervised concept learning model using a transferred concept bank, and how it can be combined with Weight of Evidence (WoE) [59].
Formally, let $f:\mathcal{X}\to\mathbb{R}^{d}$ be the pre-trained backbone model (e.g. ResNet, ResNeXt) where $\mathcal{X}$ is the input space and $d$ is the size of the embedding space. The dimension $d$ of the feature embedding space is often large (e.g., 2048 features). We then aim to reduce the dimensionality by using different techniques, either a reducer (for unsupervised learning) or a concept bank (for supervised learning). Next, we get a set of selected concepts and put them into a classifier layer $g:\mathbb{R}^{N_{c}}\to Y$ where $Y$ is the set of output classes, $N_{c}$ is the number of concepts. To implement the combined models ICE+WoE and PCBM+WoE, we replace the classifier layer of ICE and PCBM with the WoE model.
The following sections will describe in detail unsupervised concept learning and supervised concept learning based on their main difference after the feature embedding layer. For unsupervised concept learning, we use a reducer to reduce the dimensionality of the feature embedding space. The learned concepts may not be related to the concepts used in the domain; for example, they could be invalid dermatological concepts or unknown to domain experts. In contrast, the supervised concept learning technique learns the concept bank and transfers it to create a concept embedding layer. This technique requires labelling concepts on the training images, which is expensive.
Invertible Concept-based Explanation (ICE)
ICE [89] is unsupervised concept-based explanation approach that applies a reducer $R$ such as Non-negative Matrix Factorization (NMF) to reduce the dimensionality of the feature embedding space. NMF is a matrix factorization method that factorizes the feature matrix into two or more non-negative matrices [65]. We can also apply other dimensionality reduction methods such as PCA (Principal Component Analysis), which is a linear transformation method that projects the feature matrix into a lower-dimensional space. Eventually, we get a set of concepts $f_{R}(x)\in\mathbb{R}^{N_{c}}$ at the reducer layer.
Post-hoc Concept Bottleneck Model (PCBM)
PCBM [88] is a supervised concept-based explanation approach that learns the concept bank $\mathcal{C}$ and transfers it to create a concept embedding layer. We learn the concept bank by training a linear SVM to separate positive concept examples (contain the concept) and negative concept examples (do not contain the concept) in the embeddings based on CAV (Concept Activation Vector) approach [35]. Importantly, the dataset used to learn the concept bank can be different from the dataset used in the task prediction. Therefore, when the training dataset does not have concept labels, we can annotate concepts by using another dataset that has concept labels and is in the same domain.
Formally, the concept library is defined as $I=\{i_{1},i_{2},...,i_{N_{c}}\}$. The concept library can be selected by domain experts or learned from the data [25]. For each concept $i$, there are a set of feature embeddings for positive examples $P_{i}=\{f(x^{i,pos}_{1}),...,f(x^{i,pos}_{N_{pos}})\}$ and for negative examples $N_{i}=\{f(x^{i,neg}_{1}),...,f(x^{i,neg}_{N_{neg})}\}$ where $N_{pos}$ and $N_{neg}$ are the number of positive and negative examples, respectively. Next, a linear SVM is trained using $P_{i}$ and $N_{i}$ to learn CAV (the normal to the SVM’s decision boundary) for concept $i$, denoted as $c_{i}$. Finally, we get the concept matrix $\mathcal{C}\in\mathbb{R}^{N_{c}\times d}$ at the concept embedding layer.
Let $g:\mathbb{R}^{N_{c}}\to Y$ be the classifier. To learn the classifier in PCBM, we minimise the loss function:
$\min_{g}\underset{(x,y)\sim\mathcal{D}}{\mathbb{E}}[\mathcal{L}(g(f_{\mathcal{C}}(x)),y)]+\frac{\lambda}{N_{c}K}\Omega(g)$ (9)
where $f_{\mathcal{C}}(x)$ is the projection onto the concept subspace, $\mathcal{L}(\hat{y},y)$ is a loss function such as cross-entropy loss, $\Omega(g)$ is a complexity measure to regularize the model, and $\lambda$ is the regularization strength. In PCBM, a linear classifier such as a stochastic gradient descent model is implemented in this layer.
Weight of Evidence
We replace the original classifier layer in ICE and PCBM with the WoE model. Similar to the WoE model in Section 3.1, we now calculate the weight of evidence for each concept. For hypothesis $h$ and concept $c_{i}$, which is equivalent to feature $x_{i}$ in 3.1, the weight of evidence $\operatorname*{woe}$ is defined as follows.
$\operatorname*{woe}(h\mid c_{i})=\log\frac{P(c_{i}\mid h)}{P(c_{i}\mid Y_{-h})}=\log P(c_{i}\mid h)-\log P(c_{i}\mid Y_{-h})$ (10)
In the implementation of Visual Evaluative AI, we use the WoE without independence assumption.
4 Study 1: Evaluative AI on Tabular Housing Price Data
@@
In this section, we will describe the implementation of Evaluative AI on tabular data (housing price prediction domain). Experimenting with this task, we compared the hypothesis-driven approach with two state-of-the-art decision-making approaches (recommendation-driven and AI-explanation-only) using quantitative measures for efficiency, performance and reliance and qualitative analysis of information use. In the terminology of a recent review of XAI evaluation [47], the first two points of comparison are a form of evaluation with respect to the decision task, and the latter two focus on users’ perception and use of the AI system itself.
In selecting a decision-making task, we identified requirements similar to those used in other studies of how explanations can assist human decision-makers interacting with AI decision support, e.g. Vasconcelos et al. [81]: the task should not be too easy for humans to complete without a decision aid, but also, as we were using lay subjects from Prolific for this particular study, the task cannot require specialist knowledge.
We chose a version of the housing price prediction task studied previously in an XAI context [1, 18]. In this task, participants are provided with information about house features and other information which varies by experimental condition and are asked to choose whether the given house would have a sale price of low, medium or high. As noted by others [18], real estate valuation is a domain where ML models have been developed to help people make better decisions, predicting house prices is a task that lay people may need to do in real life, so it is not unrealistic to expect they have sufficient day-to-day knowledge to make predictions and decide whether or not to rely on an AI model.
@@ 4.1 Dataset and Model Implementation
We selected six features for the human experiment in the house-price decision-making task by applying a Gradient Boosting Classification model over the data. Considering domain-specific decision-making about house prices, we propose there to be three important features (quality of construction, house age and location) and three unimportant features (fireplaces, kitchen quality and central air conditioning). We divided the dataset into 80% for the training set and 20% for the test set.
@@ 4.2 Experimental Conditions
All participants [^2] We received ethics approval from our institution before conducting the human experiment (ID: 23208). were given the six house feature values plus other information, which varied by condition as set out below. Participants then chose whether the given house would have a price of low, medium or high. Using a between-subject design, participants were randomly assigned to one of three conditions:
@@ 4.3 Research Questions and Hypotheses
  • RQ4: (Information use) How do people make decisions differently in recommendation-driven, AI-explanation-only and hypothesis-driven approach?
@@
  • H1a/b: (C3) Hypothesis-driven approach will cost less time to finish the task than (C1) Recommendation-driven and (C2) AI-explanation-only.
@@
  • H2a/b: (C3) Hypothesis-driven approach will help participants make better decisions than (C1) Recommendation-driven and (C2) AI-explanation-only.
@@
For RQ4, we looked into the text written by participants when they explained why they selected an option after each question to know how they used the provided information in each decision-making approach to make their decisions. Therefore, we can identify the limitations of each approach and the generated evidence that led the participants to make a wrong decision.
@@ 4.4 Measures
    where: $C_{p,i}$ is the likelihood level of participant $p$ in question $i$, ranging from 0 to 1 (the likelihood level of a participant refers to how confident they are when answering the question); $A_{p,i}$ is the answer score of participant $p$ in question $i$, either 0 (wrong answer) or 1 (right answer); N is the number of questions for each participant. The best Brier score (i.e. equal to 0) for an individual task is when a participant answers the task correctly and gives it a 100% likelihood (or alternatively, a wrong answer but with 0% likelihood). Therefore, a participant has better task performance when they have a lower Brier score. The Brier score measures decision accuracy but awards a higher score for a correct answer when a participant is confident, and a lower penalty for an incorrect answer when not confident. This mitigates problems where participants guess answers (i.e. have low confidence in their answers). Brier score has potential for real-world applications, as it accounts for both the correctness of a decision and the confidence of the decision-maker.